# Learning Generalizable fMRI Representations for Neurological Disorder Identification

### v1

@@ Abstract
Despite the impressive advancements achieved using deep-learning for functional brain activity analysis, the heterogeneity of functional patterns and scarcity of imaging data still pose challenges in tasks such as identifying neurological disorders. For functional Magnetic Resonance Imaging (fMRI), while data may be abundantly available from healthy controls, clinical data is often scarce, especially for rare diseases, limiting the ability of models to identify clinically-relevant features. We overcome this limitation by introducing a novel representation learning strategy integrating meta-learning with self-supervised learning to improve the generalization from normal to clinical features. This approach enables generalization to challenging clinical tasks featuring scarce training data. We achieve this by leveraging self-supervised learning on the control dataset to focus on inherent features that are not limited to a particular supervised task and incorporating meta-learning to improve the generalization across domains. To explore the generalizability of the learned representations to unseen clinical applications, we apply the model to four distinct clinical datasets featuring scarce and heterogeneous data for neurological disorder classification. Results demonstrate the superiority of our representation learning strategy on diverse clinically-relevant tasks.
@@ 1 Introduction
Deep learning based approaches have demonstrated success in analyzing brain connectivity based on functional magnetic resonance imaging (fMRI) (Gadgil et al., 2020; Ahmedt-Aristizabal et al., 2021), but the scarcity and heterogeneity of fMRI data still pose challenges in clinical applications such as identifying neurological disorders. fMRI plays a vital role in identifying biomarkers for neurological disorders (Pitsik et al., 2023; Li et al., 2021). However, clinical fMRI datasets are not only high-dimensional and spatiotemporally complex but are also characterized by high variability among subjects and a limited number of data, posing significant challenges for training deep learning models to predict neurological disorders (Akrami et al., 2021). Medical datasets typically contain an abundance of healthy control data but often face a scarcity of clinical data collected for any particular neurological disorder. Simply train a model by aggregating all healthy control and clinical data may cause limited generalization and bias due to the data imbalance and heterogeneity in clinical features. This can in-turn lead to poor performance in the group with clinical pathology (Azizi et al., 2023). Moreover, deep learning models often struggle to achieve satisfactory performance on small-scale clinical datasets, frequently under-performing compared to traditional machine learning methods (Akrami et al., 2024; 2021). Considering the limitations of both data and existing deep learning models, we explore representation learning on fMRIs, aiming to extract meaningful and generalizable features from data by learning inherent functional activity patterns. Driven by the need to learn generalizable representations, we leverage self-supervised learning and meta-learning as the foundation of our representation learning approach.
@@
Our experiments are designed to demonstrate, i). the improved knowledge transfer from source to target datasets, we use a neurological disorder classification task to evaluate the performance on the target domain when applying MeTSK in a knowledge transfer task setting. ii) The generalization of representations to unseen clinical datasets. We evaluate the MeTSK model pre-trained with a source and a target dataset on unseen clinical datasets. Direct training of deep learning models on these challenging clinical datasets performed poorly due to limited training data and the heterogeneity of features, also resulting in worse performance compared to simpler machine learning classifiers. So acquiring generalizable representations is crucial for these clinical datasets. According to (Kumar et al., 2022), fine-tuning can distort good pre-trained features and degrade downstream performance under large distribution shifts. Here we explored linear probing for evaluating the generalization of representations on distinct neurological disorder classification tasks. As we show below, we are able to consistently achieve superior classification performance for diverse neurological disorder identification tasks compared to linear classification directly using traditional functional connectivity features as input.
@@ 2 Related Work
In fMRI analysis, heterogeneity and disparities across datasets pose significant challenges for the generalization of models. Most existing methods developed for fMRIs focus on adapting models between closely related domains, such as generalizing across imaging sites within the same dataset. These methods are often designed to address site-specific variations, such as differences in imaging protocols or scanner types (Li et al., 2020; Shi et al., 2021; Liu et al., 2023). Other approaches, such as fine-tuning (Raghu et al., 2019) and multi-task learning (Huang et al., 2020), attempt to adapt to target domains but often fail to learn transferable features due to domain discrepancies (Liu et al., 2020b; Raghu et al., 2019) and may distort the learning of target-domain specific features Chen et al. (2019).
@@ 3 Methods
In this section, we introduce our proposed strategy, MeTSK, which improves the generalization of self-supervised fMRI features from a control dataset to clinical datasets. The proposed network architecture consists of a feature extractor that learns generalizable features from both source (control) and target (clinical) domains, and source and target heads to learn domain-specific features for the source and target domain, respectively. The bi-level optimization strategy is applied to learn generalizable features using a Spatio-temporal Graph Convolutional Network (ST-GCN) (Gadgil et al., 2020) as the backbone model. After MeTSK is trained on the control and clinical datasets, its generalization capability will be evaluated on unseen clinical datasets using linear probing. The methodology of MeTSK as well as the representation learning pipeline are shown in Fig. 1.
@@ 3.1 Feature Extractor: ST-GCN
We adopt a popular model for fMRI classification, ST-GCN (Gadgil et al., 2020), as the backbone architecture to extract graph representations from both spatial and temporal information. A graph convolution and a temporal convolution are performed in one ST-GCN module shown in Fig. 2, following the details in (Gadgil et al., 2020). The feature extractor includes three ST-GCN modules. The target head and the source head share the same architecture, which consists of one ST-GCN module and one fully-connected layer.
@@
Figure 2: An illustration of the ST-GCN model architecture. Spatial graph convolution is first applied to the spatial graph at each time point. Then 1D temporal convolutions are performed along the resulting features on each node. Multiple sub-sequences are randomly sampled from the whole time series as input graphs for training.
@@ 3.3 Contrastive Self-supervised Learning
The optimization objective $\mathcal{L_{T}}$ of the target domain depends on the target task. In a classification task with class labels $Y_{\mathcal{T}}$, we adopt the Cross-Entropy loss. The total loss for the proposed MeTSK strategy is then:
@@ 3.4 Linear Probing
The bi-level optimization strategy and contrastive self-supervised learning work together to learn robust, domain-invariant features, enhancing the ability to generalize effectively to unseen clinical datasets. Here we evaluate the generalization of learned representations through linear probing, a crucial method for evaluating the quality of representations learned by the model (Chen et al., 2020a; Kumar et al., 2022). Linear probing involves freezing the parameters of a pre-trained model and training a linear classifier on the output. The intuition behind linear probing is that good representations should be linearly separable between classes (Chen et al., 2020a). We apply the MeTSK model pre-trained on the source and target domains to directly generate features for the unseen fMRI data without any fine-tuning. We then input these features into a linear classifier to perform neurological disorder classification.
@@ 4 Datasets
We use HCP (Van Essen et al., 2013) as our source dataset due to its large size. For target datasets, we use the ADHD (Bellec et al., 2017) datasets, and the ABIDE dataset (Craddock et al., 2013) during the training of the MeTSK model. We then introduce four independent clinical datasets as held-out domains for evaluating generalization performance.
@@ 4.1 HCP Dataset
The healthy control data (source domain) is drawn from the Human Connectome Project (HCP) S1200 dataset (Van Essen et al., 2013). The HCP database includes 1,096 young adult (ages 22-35) subjects with resting-state-fMRI data collected at a total of 1200 time-points for each of four sessions. The preprocessing of fMRI follows the minimal preprocessing procedure in (Gadgil et al., 2020; Glasser et al., 2013). Finally, the brain was parcellated into $116$ Regions of Interest (ROIs) using the Automated Anatomical Labeling (AAL) atlas in (Tzourio-Mazoyer et al., 2002). The AAL atlas was defined based on brain anatomy. It divides the brain into 116 regions, including 90 cerebrum regions and 26 cerebellum regions. These 116 regions form the nodes of our graph. The fMRI data were reduced to a single time-series per node by averaging across each ROI.
@@ 4.2 ADHD-Peking Dataset
The Attention-Deficit/Hyperactivity Disorder (ADHD-200) consortium data from the Peking site (Bellec et al., 2017) includes 245 subjects in total, with 102 ADHD subjects and 143 Typically Developed Controls (TDC). To model the situation where the clinical target data set is small, we use only the subset of the larger ADHD database that was collected from the Peking site. In the ablation studies, we explore the impact of using the complete set. We use the preprocessed data released on (http://preprocessed-connectomes-project.org/adhd200/). During preprocessing, the initial steps involve discarding the first four time points, followed by slice time and motion correction. The data is then registered to the Montreal Neurological Institute (MNI) space, processed with a band-pass filter (0.009Hz - 0.08Hz), and smoothed using a 6 mm Full Width at Half Maximum (FWHM) Gaussian filter. The fMRI data consisted of $231$ time points after preprocessing. As a final step, the ADHD-Peking data were re-registered from MNI space to the same AAL atlas as for the HCP subjects, and the average time-series computed for each ROI.
@@ 4.3 ABIDE-UM Dataset
The Autism Brain Imaging Data Exchange I (ABIDE I) (Craddock et al., 2013) collects resting-state fMRI from 17 international sites. Similar to the ADHD dataset, we use only the subset of data from the UM site, which includes 66 subjects with Autism Spectrum Disorder (ASD) and 74 TDCs (113 males and 27 females aged between 8-29). We downloaded the data from http://preprocessed-connectomes-project.org/abide/, where data was pre-processed using the C-PAC pre-processing pipeline (Craddock et al., 2013). The fMRI data underwent several preprocessing steps: slice time correction, motion correction, and voxel intensity normalization. The data was then band-pass filtered (0.01–0.1 Hz) and spatially registered to the MNI152 template space using a nonlinear method. All fMRIs have 296 time points. As a final step, the ABIDE-UM data were re-registered from MNI space to the same AAL atlas as for the HCP subjects, and the average time-series computed for each ROI.
@@ 4.4 Post-traumatic Epilepsy Dataset
Pre-processing of the MagNeTs rs-fMRI data was performed using the BrainSuite fMRI Pipeline (BFP) (https://brainsuite.org). BFP is a software workflow that processes fMRI and T1-weighted MR data using a combination of software that includes BrainSuite, AFNI, FSL, and MATLAB scripts to produce processed fMRI data represented in a common grayordinate system that contains both cortical surface vertices and subcortical volume voxels (Glasser et al., 2013). As described above, the pre-processed data were then mapped to the same AAL atlas as used with the other datasets. Regional time-series were then generated for each of the 116 parcels by averaging over the corresponding region of interest.
@@ 4.5 Alzheimer’s Disease Dataset
Open Access Series of Imaging Studies (OASIS-3) (LaMontagne et al., 2019) is a longitudinal neuroimaging, clinical, and cognitive dataset for normal aging and Alzheimer Disease (AD), including 1379 participants: 755 cognitively normal adults and 622 individuals at various stages of cognitive decline ranging in age from 42-95 years old. Resting-state fMRI data was used for the classification of Alzheimer’s Disease and is publicly available at https://www.oasis-brains.org. We randomly selected a subset of the OASIS-3 dataset, consisting of 42 subjects diagnosed as AD and 42 cognitively normal subjects for the downstream clinical task. We are using a small subset to simulate the case where only limited disease data is available. There are 164 time points in each fMRI scan. The preprocessing steps used for OASIS-3 dataset are the same as used for the PTE dataset. We perform binary classification between PD and control subjects.
@@ 4.6 Parkinson’s Disease Datasets
Neurocon Dataset: The Neurocon dataset is provided by the Neurology Department of the University Emergency Hospital Bucharest (Romania) Badea et al. (2017), which includes 27 PD patients and 16 normal controls (with 2 replicate scans per subject). Both the rs-fMRI and T1-weighted scans were collected from a 1.5-Tesla Siemens Avanto MRI scanner. All fMRI scans have 137 time points. Both the Neurocon and TaoWU datasets were downloaded from https://fcon_1000.projects.nitrc.org/indi/retro/parkinsons.html. We pre-processed TaoWu and Neurocon fMRI data using the same preprocessing pipeline (BFP) that was applied to the PTE dataset.
@@ 5.1 Evaluation of Representation Transferability
For comparison, we designed (i) a baseline model using a ST-GCN with a supervised task directly trained on the target dataset (Baseline), (ii) a ST-GCN model fine-tuned on the target dataset after pre-training on HCP data (FT), (iii) a model performing multi-task learning on both source and target datasets simultaneously (MTL), and (iv) the proposed strategy, MeTSK. We incorporated MTL and FT methods for comparison in order to investigate whether MeTSK is superior to traditional approaches in terms of knowledge transfer. We compared several baseline methods: a Linear Support Vector Machine (SVM), a Random Forest Classifier (RF), a Multi-Layer Perceptron (MLP) consisting of three linear layers, an LSTM model for fMRI analysis (Gadgil et al., 2020), and a model combining a transformer and graph neural network (STAGIN) (Kim et al., 2021). For the SVM, RF, and MLP, the inputs are flattened functional connectivity features, calculated using the Pearson’s correlation coefficient between fMRI time-series across pairs of brain regions defined in the AAL atlas. LSTM and STAGIN, on the other hand, utilize raw fMRI time-series as their input.

### v2

# Generalizable Representation Learning for fMRI-based Neurological Disorder Identification

@@ Abstract
Despite the impressive advances achieved using deep learning for functional brain activity analysis, the heterogeneity of functional patterns and the scarcity of imaging data still pose challenges in tasks such as identifying neurological disorders. For functional Magnetic Resonance Imaging (fMRI), while data may be abundantly available from healthy controls, clinical data is often scarce, especially for rare diseases, limiting the ability of models to identify clinically-relevant features. We overcome this limitation by introducing a novel representation learning strategy integrating meta-learning with self-supervised learning to improve the generalization from normal to clinical features. This approach enables generalization to challenging clinical tasks featuring scarce training data. We achieve this by leveraging self-supervised learning on the control dataset to focus on inherent features that are not limited to a particular supervised task and incorporating meta-learning to improve the generalization across domains. To explore the generalizability of the learned representations to unseen clinical applications, we apply the model to four distinct clinical datasets featuring scarce and heterogeneous data for neurological disorder classification. Results demonstrate the superiority of our representation learning strategy on diverse clinically-relevant tasks. Code is publicly available at https://github.com/wenhui0206/MeTSK/tree/main
@@ 1 Introduction
Deep learning based approaches have demonstrated success in analyzing brain connectivity based on functional magnetic resonance imaging (fMRI) (Gadgil et al., 2020; Ahmedt-Aristizabal et al., 2021), but the scarcity and heterogeneity of fMRI data still pose challenges in clinical applications such as identifying neurological disorders. fMRI plays a vital role in identifying biomarkers for neurological disorders (Pitsik et al., 2023; Li et al., 2021). However, clinical fMRI datasets are not only high-dimensional and spatiotemporally complex but are also characterized by high variability among subjects and a limited number of data, posing significant challenges for training deep learning models to predict neurological disorders (Akrami et al., 2021). Medical datasets typically contain an abundance of healthy control data but often face a scarcity of clinical data collected for any particular neurological disorder. Simply training a model by aggregating all healthy control and clinical data may cause limited generalization and bias due to the data imbalance and heterogeneity in clinical features. This can in-turn lead to poor performance in the group with clinical pathology (Azizi et al., 2023). Moreover, deep learning models often struggle to achieve satisfactory performance on small-scale clinical datasets, frequently under-performing compared to traditional machine learning methods (Akrami et al., 2024; 2021). Considering the limitations of both data and existing deep learning models, we explore representation learning on fMRIs, aiming to extract meaningful and generalizable features from data by learning inherent functional activity patterns. Driven by the need to learn generalizable representations, we leverage self-supervised learning and meta-learning as the foundation of our representation learning approach.
@@
Our experiments are designed to demonstrate, i). the improved knowledge transfer from source to target datasets, we use a neurological disorder classification task to evaluate the performance on the target domain when applying MeTSK in a knowledge transfer task setting. ii) The generalization of representations to unseen clinical datasets. We evaluate the MeTSK model pre-trained with a source and a target dataset on unseen clinical datasets. Direct training of deep learning models on these challenging clinical datasets performed poorly due to limited training data and the heterogeneity of features, also resulting in worse performance compared to simpler machine learning classifiers. So acquiring generalizable representations is crucial for these clinical datasets. According to Kumar et al. (2022), fine-tuning can distort good pre-trained features and degrade downstream performance under large distribution shifts. Here we explored linear probing for evaluating the generalization of representations on distinct neurological disorder classification tasks. As we show below, we are able to consistently achieve superior classification performance for diverse neurological disorder identification tasks compared to linear classification directly using traditional functional connectivity features as input.
@@ 2 Related Work
In fMRI analysis, heterogeneity and disparities across datasets pose significant challenges for the generalization of models. Most existing methods developed for fMRIs focus on adapting models between closely related domains, such as generalizing across imaging sites within the same dataset. These methods are often designed to address site-specific variations, such as differences in imaging protocols or scanner types (Li et al., 2020; Shi et al., 2021; Liu et al., 2023). Other approaches, such as fine-tuning (Raghu et al., 2019) and multi-task learning (Huang et al., 2020), attempt to adapt to target domains but often fail to learn transferable features due to domain discrepancies (Liu et al., 2020b; Raghu et al., 2019) and may distort the learning of target-domain specific features (Chen et al., 2019).
@@ 3 Methods
In this section, we introduce our proposed strategy, MeTSK, which improves the generalization of self-supervised fMRI features from a control dataset to clinical datasets. The proposed network architecture consists of a feature extractor that learns generalizable features from both source (control) and target (clinical) domains, and source and target heads to learn domain-specific features for the source and target domain, respectively. The bi-level optimization strategy is applied to learn generalizable features using a Spatio-temporal Graph Convolutional Network (ST-GCN) (Gadgil et al., 2020) as the backbone model. After MeTSK is trained on the control and clinical datasets, its generalization capability will be evaluated on unseen clinical datasets using linear probing. The methodology of MeTSK as well as the representation learning pipeline are shown in Figure 1.
@@ 3.1 Feature Extractor: ST-GCN
We adopt a popular model for fMRI classification, ST-GCN (Gadgil et al., 2020), as the backbone architecture to extract graph representations from both spatial and temporal information. A graph convolution and a temporal convolution are performed in one ST-GCN module shown in Fig. 2, following the details in Gadgil et al. (2020). The feature extractor includes three ST-GCN modules. The target head and the source head share the same architecture, which consists of one ST-GCN module and one fully-connected layer.
@@
Figure 2: An illustration of the ST-GCN model architecture. Spatial graph convolution is first applied to the spatial graph at each time point. Then 1D temporal convolutions are performed along the resulting features on each node. Multiple sub-sequences are randomly sampled from the whole time series as input graphs.
@@ 3.3 Contrastive Self-supervised Learning
The optimization objective $\mathcal{L_{T}}$ of the target domain depends on the target task. In a classification task with class labels $Y_{\mathcal{T}}$, we adopt the Cross-Entropy loss. The total loss for the proposed MeTSK strategy is:
@@ 3.4 Linear Probing
The bi-level optimization strategy and contrastive self-supervised learning work together to learn robust, domain-invariant features, enhancing the ability to generalize effectively to unseen clinical datasets. Here we evaluate the generalization of learned representations through linear probing, a crucial method for evaluating the quality of representations learned by the model (Chen et al., 2020a; Kumar et al., 2022). Linear probing involves freezing the parameters of a pre-trained model and training a linear classifier on the output. The intuition behind linear probing is that good representations should be linearly separable between classes (Chen et al., 2020a). We apply the MeTSK model pre-trained on the source and target domains to directly generate features for the unseen fMRI data without any fine-tuning. We then input these features into a linear classifier to perform neurological disorder classification. In alignment with common practices in the literature (Ortega Caro et al., 2023; He et al., 2020; Grill et al., 2020), we adopted linear SVM and logistic regression as probing classifiers. Both produce linear decision boundaries, enabling evaluation of representation quality through linear separability.
@@ 4 Datasets
We use HCP (Van Essen et al., 2013) as our source dataset due to its large size. For target datasets, we use the ADHD (Bellec et al., 2017) datasets, and the ABIDE dataset (Craddock et al., 2013) during the training of the MeTSK model. We then introduce four independent clinical datasets as held-out domains for evaluating generalization performance. For all the fMRI datasets used in this work, we parcellate the brain into $116$ Regions of Interest (ROIs) using the Automated Anatomical Labeling (AAL) atlas in Tzourio-Mazoyer et al. (2002). The AAL atlas was defined based on brain anatomy. It divides the brain into 116 regions, including 90 cerebrum regions and 26 cerebellum regions. These 116 regions form the nodes of our graph. The fMRI data were reduced to a single time series per node by averaging the voxel-level signals within each region of interest (ROI). We then computed pairwise Pearson correlations between ROIs to obtain a functional connectivity matrix of size $116\times 116$. The upper triangular part, including the diagonal, was extracted and flattened to form the input connectivity features for machine learning classifiers.
@@ 4.1 HCP Dataset
The healthy control data (source domain) is drawn from the Human Connectome Project (HCP) S1200 dataset (Van Essen et al., 2013). The HCP database includes 1,096 young adult (ages 22-35) subjects with resting-state-fMRI data collected at a total of 1200 time-points for each of four sessions. We used publicly available preprocessed fMRI data that follow the minimal preprocessing pipelines described in Glasser et al. (2013). Pre-processing steps include spatial distortion correction using field maps, motion correction through rigid-body registration, and EPI distortion correction. Functional images are registered to each subject’s structural T1-weighted image, followed by spatial normalization to Montreal Neurological Institute (MNI) space. The data are then processed to align cortical data to a standard surface space (grayordinate system).
@@ 4.2 ADHD-Peking Dataset
The Attention-Deficit/Hyperactivity Disorder (ADHD-200) consortium data from the Peking site (Bellec et al., 2017) includes 245 subjects in total, with 102 ADHD subjects and 143 Typically Developed Controls (TDC). To model the situation where the clinical target data set is small, we use only the subset of the larger ADHD database that was collected from the Peking site. In the ablation studies, we explore the impact of using the complete set. We use the preprocessed data released on (http://preprocessed-connectomes-project.org/adhd200/). During preprocessing, the initial steps involve discarding the first four time points, followed by slice time and motion correction. The data is then registered to the MNI space, processed with a band-pass filter (0.009Hz - 0.08Hz), and smoothed using a 6 mm Full Width at Half Maximum (FWHM) Gaussian filter. All fMRIs have $231$ time points after preprocessing.
@@ 4.3 ABIDE-UM Dataset
The Autism Brain Imaging Data Exchange I (ABIDE I) (Craddock et al., 2013) collects resting-state fMRI from 17 international sites. Similar to the ADHD dataset, we use only the subset of data from the UM site, which includes 66 subjects with Autism Spectrum Disorder (ASD) and 74 TDCs (113 males and 27 females aged between 8-29). We downloaded the data from http://preprocessed-connectomes-project.org/abide/, where data was pre-processed using the C-PAC pre-processing pipeline (Craddock et al., 2013). The fMRI data underwent several preprocessing steps: slice time correction, motion correction, and voxel intensity normalization. The data was then band-pass filtered (0.01–0.1 Hz) and spatially registered to the MNI template space using a nonlinear method. Spatial smoothing was applied using a 6mm FWHM Gaussian kernel to enhance the signal-to-noise ratio. All fMRIs have 296 time points.
@@ 4.4 Post-traumatic Epilepsy Dataset
Pre-processing of the MagNeTs rs-fMRI data was performed using the BrainSuite fMRI Pipeline (BFP) (https://brainsuite.org). BFP is a software workflow that processes fMRI and T1-weighted MR data using a combination of software that includes BrainSuite, AFNI, FSL, and MATLAB scripts to produce processed fMRI data represented in a common grayordinate system that contains both cortical surface vertices and subcortical volume voxels (Glasser et al., 2013). The pre-processing follow the same minimal processing pipeline used for HCP data.
@@ 4.5 Alzheimer’s Disease Dataset
Open Access Series of Imaging Studies (OASIS-3) (LaMontagne et al., 2019) is a longitudinal neuroimaging, clinical, and cognitive dataset for normal aging and Alzheimer Disease (AD), including 1379 participants: 755 cognitively normal adults and 622 individuals at various stages of cognitive decline ranging in age from 42-95 years old. Resting-state fMRI data was used for the classification of Alzheimer’s Disease and is publicly available at https://www.oasis-brains.org. We randomly selected a subset of the OASIS-3 dataset, consisting of 42 subjects diagnosed as AD and 42 cognitively normal subjects for the downstream clinical task. We are using a small subset to simulate the case where only limited disease data is available. There are 164 time points in each fMRI scan. The preprocessing steps used for OASIS-3 dataset are the same as used for the PTE dataset. We perform binary classification between AD and control subjects.
@@ 4.6 Parkinson’s Disease Datasets
Neurocon Dataset: The Neurocon dataset is provided by the Neurology Department of the University Emergency Hospital Bucharest (Romania) (Badea et al., 2017), which includes 27 PD patients and 16 normal controls (with 2 replicate scans per subject). Both the rs-fMRI and T1-weighted scans were collected from a 1.5-Tesla Siemens Avanto MRI scanner. All fMRI scans have 137 time points. Both the Neurocon and TaoWU datasets were downloaded from https://fcon_1000.projects.nitrc.org/indi/retro/parkinsons.html. We pre-processed TaoWu and Neurocon fMRI data using the same preprocessing pipeline (BFP) that was applied to the PTE dataset.
@@ 5.1 Evaluation of Representation Transferability
For comparison, we designed (i) a baseline model using a ST-GCN with a supervised task directly trained on the target dataset (Baseline), (ii) a ST-GCN model fine-tuned on the target dataset after pre-training on HCP data (FT), (iii) a model performing multi-task learning on both source and target datasets simultaneously (MTL), and (iv) the proposed strategy, MeTSK. We incorporated MTL and FT methods for comparison in order to investigate whether MeTSK is superior to traditional approaches in terms of knowledge transfer. We compared several baseline methods: a Linear Support Vector Machine (SVM), a Random Forest Classifier (RF), a Multi-Layer Perceptron (MLP) consisting of three linear layers, an LSTM model for fMRI analysis (Gadgil et al., 2020), and a model combining a transformer and graph neural network (STAGIN) (Kim et al., 2021). For the SVM, RF, and MLP, the inputs are flattened functional connectivity features, calculated using the Pearson’s correlation coefficient between fMRI time-series across pairs of brain regions defined in the AAL atlas. LSTM and STAGIN use raw fMRI time-series as their input.